# thanks: Author to whom correspondence should be addressed:

### v1

@@ I Introduction
Neural-network Quantum States (NQS)Carleo and Troyer (2017) have emerged as a promising avenue, utilizing artificial neural networks to encapsulate the many-body wave function in a compact, second-quantized formulation. This innovative ansatz has shown remarkable successHermann et al. (2023) in molecularChoo, Mezzacapo, and Carleo (2020); Yang et al. (2020a); Barrett, Malyshev, and Lvovsky (2022); Zhao, Stokes, and Veerapaneni (2023); Li et al. (2023); Wu et al. (2023); Liu and Clark (2024) and solid-stateYoshioka, Mizukami, and Nori (2021) systems and has been extended to continuous spaces.Hermann, Schätzle, and Noé (2020); Pfau et al. (2020); von Glehn, Spencer, and Pfau (2022); Li et al. (2022); Ren et al. (2023); Li et al. (2024) The energy evaluation and NQS ansatz optimization usually require the Orbital Space Variational Monte Carlo (VMC) framework, whose typical complexity is $O(N^{5})$ but can be improved to $O(N^{4})$Wei and Neuscamman (2018) or lowerSabzevari and Sharma (2018) with strategic approximations, where $N$ is the system size. Nonetheless, the computational overhead introduced by neural networks, compared to traditional wave function ansatzes, has limited NQS’s broader applications in molecules. A notable challenge lies in the inefficiency of Monte Carlo sampling within the VMC framework. One of the promising solutions is exact sampling via autoregressive neural networks,Barrett, Malyshev, and Lvovsky (2022) as demonstrated with the weakly correlated $\rm Na_{2}CO_{3}$ in minimal basis sets, showcasing competitive accuracy to the Coupled Cluster with Singles and Doubles (CCSD) method.Zhao, Stokes, and Veerapaneni (2023) More recently, we introduced another deterministic configuration selection scheme to avoid Monte Carlo sampling,Li et al. (2023) eliminating stochastic noise and enhancing the applicability of various neural network types.Liu and Clark (2024) Building on this insight, in this article, we detail three algorithmic advancements to optimize NQS through orbital space VMC calculations and examine the performance of NQS predicated on Restricted Boltzmann Machine (RBM) model in tackling large, strongly correlated molecular systems.
@@ II.1 Overview
where $T=1$ is an hyperparameter, and $\epsilon_{i}$ (or $a_{i}$) are updated during NQS optimization. The Boltzmann distribution can be considered as an independent electron approximation, applicable to various NQS models.
@@ II.4 Block optimization for the restricted Boltzmann machine
This decomposition illustrates how the SR method can be applied to each block individually, thus facilitating the optimization of significantly large RBM models. The strategy of dividing parameters into blocks can be static, remaining unchanged throughout the optimization process, or dynamic, with blocks being re-divided at each parameter update. Furthermore, our experiments have demonstrated the efficacy of optimizing a subset of the blocked parameters at a time. To aid visualization, we include a graphic representation (Figure 2) of the RBM division when $N_{b}=2$, underscoring the practical implementation of this block optimization strategy.
@@ III.1 Performance of adaptive learning rate
As illustrated in Figure 3, both adaptive and constant learning rate approaches achieve comparable levels of high accuracy. However, it is noteworthy that the relatively large constant learning rates (0.2, 0.25 and 0.30) result in some instances of failure. Regarding efficiency, as depicted in Figure 4, the adaptive learning rate scheme outperforms the constant rate in terms of both wall time and iteration count to achieve similar accuracy levels when $\eta_{0}>0.05$. This efficiency gain stems from the adaptive method’s capability to judiciously utilize small learning rates for stability and large rates for faster convergence, avoiding the oscillations near the minimum often observed with larger constant rates. This is corroborated by the observation that the large constant learning rates exhibit a higher incidence of non-convergence within 10,000 iterations. Interestingly, in cases where $\eta_{0}\leq 0.05$, the adaptive learning rate algorithm consistently selects the maximum available value ($\eta_{0}$), aligning with the behavior of the constant learning rate scheme in terms of iteration count. Besides,the distinctions in wall time between adaptive and constant methods reveal that the adaptive scheme incurs merely about twice the time cost per iteration due to its additional energy evaluation, a reasonable trade-off for its considerable improvements in learning rate selection at larger $\eta_{0}$ values.
@@ III.3 Performance of block optimization
Results depicted at the top of Figure 7 illustrate that all variations of block optimization, with the exception of the scheme optimizing only a fixed half of the parameters, achieve higher accuracy than conventional optimization approaches, regardless of whether $\alpha$ is 2 or 4. Notably, a dynamic division into 2 blocks—regardless of optimizing half or all blocked parameters—consistently reaches chemical accuracy in nearly all instances. Moreover, the 2-block optimization marginally outperforms the 4-block strategy, with both fixed and dynamic approaches yielding similar results. It’s important to highlight that optimizing a fixed half of the parameters shows comparable performance to conventional optimization when $\alpha=2$, as both essentially optimize an RBM model with an equivalent parameter count.
@@ III.5 Ground-state energy of $\mathbf{Cr_{2}}$
The optimization begins with a CASSCF(12e,12o) calculation, utilizing the derived natural molecular orbitals. The adaptive learning rate algorithm is implemented, dynamically choosing from 100 potential values within the range $\left[{10}^{-4},0.2\right]$. Constrained optimization, with $N_{\rm opt}=5$, is applied to ensure stable energy reduction. For block optimization, the RBM model, with a hidden unit density of $\alpha=12$, is dynamically partitioned into three blocks at each update, optimizing parameters in two of the three blocks. The input bias of the RBM model is initiated with a hyperparameter $T=1/6$, and the diagonal regularization parameter ($\lambda$) for the SR method is $10^{-5}$. To strike a balance between the accuracy and efficiency, the selected configuration sample $\mathcal{V}$ includes electron configurations with $\left|\psi_{\theta}\right|>4\times{10}^{-5}$ for NQS optimization.
As revealed in TABLE 1, after approximately 70 optimization iterations, the NQS’s variational energies in the active space already exceed the accuracy of the CCSDT method. The total energies, after perturbative correction (detailed in Appendix A), outperform both the CCSD(T) and CCSDTQ methods.
@@
Initially selecting around 170,000 electron configurations with $\left|\psi_{\theta}\right|>1\times{10}^{-4}$ using the optimized NQS ansatz, the resultant variational energy $\widetilde{E}_{\theta}$ differs from the HCI energy with around 120,000 electron configurations by approximately 4mHa, with a perturbatively corrected energy discrepancy of about 3mHa. This suggests a high quality of the current truncated NQS ansatz close to the HCI wave function. A subsequent selection of roughly 390,000 electron configurations with $\left|\psi_{\theta}\right|>4\times{10}^{-5}$ yields a minor decrease (less than 4mHa) in variational energy, contrasting with a significant drop (about 15mHa) in the variational HCI energy with about 480,000 configurations. This indicates that enhancing NQS accuracy requires a much larger configuration sample $\mathcal{V}$ during optimization.
@@
NQS 146 974 $\mathbf{-0.38761}$ $\mathbf{-0.43805}$
NQS 313 675 $\mathbf{-0.38955}$ $\mathbf{-0.43513}$
@@ Acknowledgements.
Xiang Li: Conceptualization (equal); Data Curation (equal); Formal Analysis (equal); Methodology (equal); Project Administration (supporting); Software (equal); Supervision (supporting); Validation (equal); Visualization (equal); Writing – original draft (lead); Writing – review & editing (equal) Jia-Cheng Huang: Conceptualization (equal); Data Curation (equal); Formal Analysis (equal); Methodology (equal); Software (equal); Validation (equal); Visualization (equal); Writing – review & editing (equal) Guang-Ze Zhang: Conceptualization (equal); Data Curation (equal); Formal Analysis (equal); Methodology (equal); Software (equal); Validation (equal); Visualization (equal); Writing – review & editing (equal) Hao-En Li: Validation (supporting); Visualization (supporting); Writing – review & editing (equal) Zhu-Ping Shen: Validation (supporting); Visualization (supporting); Writing – review & editing (equal) Chen Zhao: Visualization (supporting); Writing – review & editing (supporting) Han-Shi Hu: Funding Acquisition (lead); Project Administration (lead); Resources (lead); Supervision (lead); Writing – review & editing (equal)
@@
Appendix A Derivation of the perturbative correction for NQS

### v3

@@ I Introduction
Neural-network Quantum States (NQS)Carleo and Troyer (2017) have emerged as a promising avenue, utilizing artificial neural networks to encapsulate the many-body wave function in a compact, second-quantized formulation. This innovative ansatz has shown remarkable successHermann et al. (2023) in molecularChoo, Mezzacapo, and Carleo (2020); Yang et al. (2020a); Barrett, Malyshev, and Lvovsky (2022); Zhao, Stokes, and Veerapaneni (2023); Li et al. (2023); Wu et al. (2023); Liu and Clark (2024) and solid-stateYoshioka, Mizukami, and Nori (2021) systems, and has been extended to both continuous spacesHermann, Schätzle, and Noé (2020); Pfau et al. (2020); von Glehn, Spencer, and Pfau (2022); Li et al. (2022); Ren et al. (2023); Li et al. (2024) and quantum computing.Xia and Kais (2018); Hagai et al. (2023); Torlai et al. (2018, 2020); Bennewitz et al. (2022) The energy evaluation and NQS ansatz optimization usually require the Orbital Space Variational Monte Carlo (VMC) framework, whose typical complexity is $O(N^{5})$ but can be improved to $O(N^{4})$Wei and Neuscamman (2018) or lowerSabzevari and Sharma (2018) with strategic approximations, where $N$ is the system size. Nonetheless, the computational overhead introduced by neural networks, compared to traditional wave function ansatzes, has limited NQS’s broader applications in molecules. A notable challenge lies in the inefficiency of Monte Carlo sampling within the VMC framework. One of the promising solutions is exact sampling via autoregressive neural networks,Barrett, Malyshev, and Lvovsky (2022) as demonstrated with the weakly correlated $\rm Na_{2}CO_{3}$ in minimal basis sets, showcasing competitive accuracy to the Coupled Cluster with Singles and Doubles (CCSD) method.Zhao, Stokes, and Veerapaneni (2023) More recently, we introduced another deterministic configuration selection scheme to avoid Monte Carlo sampling,Li et al. (2023) eliminating stochastic noise and enhancing the applicability of various neural network types.Liu and Clark (2024) Building on this insight, in this article, we detail three algorithmic advancements to optimize NQS through orbital space VMC calculations and examine the performance of NQS predicated on Restricted Boltzmann Machine (RBM) model in tackling large, strongly correlated molecular systems.
@@ II.1 Overview
where $T=1$ is an hyperparameter, and $\epsilon_{i}$ (or $a_{i}$) are updated during NQS optimization. The Boltzmann distribution can be considered as an independent electron approximation, applicable to various NQS models, as detailed in Appendix A.
@@ II.4 Block optimization for the restricted Boltzmann machine
This decomposition illustrates how the SR method can be applied to each block individually, thus facilitating the optimization of significantly large RBM models. The strategy of dividing parameters into blocks can be static, remaining unchanged throughout the optimization process, or dynamic, with blocks being re-divided at each parameter update. Furthermore, our experiments have demonstrated the efficacy of optimizing a subset of the blocked parameters at a time. To aid visualization, we include a graphic representation (Figure 2) of the RBM division when $N_{b}=2$, underscoring the practical implementation of this block optimization strategy. We note that Zhang et al.Zhang et al. (2023) have recently also proposed one sequential local optimization strategy for RBM, which exploits the locality of the Hamiltonian and a segmentation of the input units.
@@ III.1 Performance of adaptive learning rate
As illustrated in Figure 3, both adaptive and constant learning rate approaches achieve comparable levels of high accuracy. However, it is noteworthy that the relatively large constant learning rates (0.2, 0.25 and 0.30) result in some instances of falling into poor local minima, whereas the corresponding adaptive approaches do not. Regarding efficiency, as depicted in Figure 4, the adaptive learning rate scheme outperforms the constant rate in terms of both wall time and iteration count to achieve similar accuracy levels when $\eta_{0}>0.05$. This efficiency gain stems from the adaptive method’s capability to judiciously utilize small learning rates for stability and large rates for faster convergence, avoiding the oscillations near the minimum often observed with larger constant rates. This is corroborated by the observation that the large constant learning rates exhibit a higher incidence of non-convergence within 10,000 iterations. Interestingly, in cases where $\eta_{0}\leq 0.05$, the adaptive learning rate algorithm consistently selects the maximum available value ($\eta_{0}$), aligning with the behavior of the constant learning rate scheme in terms of iteration count. Besides,the distinctions in wall time between adaptive and constant methods reveal that the adaptive scheme incurs merely about twice the time cost per iteration due to its additional energy evaluation, a reasonable trade-off for its considerable improvements in learning rate selection at larger $\eta_{0}$ values.
@@ III.3 Performance of block optimization
Results depicted at the top of Figure 7 illustrate that all variations of block optimization, with the exception of the scheme optimizing only a fixed half of the parameters, achieve higher accuracy than conventional optimization approaches, regardless of whether $\alpha$ is 2 or 4. Notably, a dynamic division into 2 blocks—regardless of optimizing half or all blocked parameters—consistently reaches chemical accuracy in nearly all instances. This demonstrates that the noise introduced by block optimization in Eq. (10) can reduce the likelihood of the NQS getting stuck in local minima. Moreover, the 2-block optimization marginally outperforms the 4-block strategy, with both fixed and dynamic approaches yielding similar results. It’s important to highlight that optimizing a fixed half of the parameters shows comparable performance to conventional optimization when $\alpha=2$, as both essentially optimize an RBM model with an equivalent parameter count.
@@ III.5 Ground-state energy of $\mathbf{Cr_{2}}$
The optimization begins with a CASSCF(12e,12o) calculation, utilizing the derived natural molecular orbitals. The adaptive learning rate is implemented, dynamically choosing from 100 potential values within the range $\left[{10}^{-4},0.2\right]$. Constrained optimization, with $N_{\rm opt}=5$, is applied to ensure stable energy reduction. For block optimization, the RBM model, with a hidden unit density of $\alpha=12$, is dynamically partitioned into three blocks at each update, optimizing parameters in two of the three blocks. The input bias of the RBM model is initiated with a hyperparameter $T=1/6$, and the diagonal regularization parameter ($\lambda$) for the SR method is $10^{-5}$. To strike a balance between the accuracy and efficiency, the selected configuration sample $\mathcal{V}$ includes electron configurations with $\left|\psi_{\theta}\right|>4\times{10}^{-5}$ for NQS optimization.
As revealed in TABLE 1, after approximately 70 optimization iterations, the NQS’s variational energies in the active space already exceed the accuracy of the CCSDT method. The total energies, after perturbative correction (detailed in Appendix B), outperform both the CCSD(T) and CCSDTQ methods. Moreover, initially selecting around 170,000 electron configurations with $\left|\psi_{\theta}\right|>1\times{10}^{-4}$ using the optimized NQS ansatz, the resultant variational energy $\widetilde{E}_{\theta}$ differs from the HCI energy with around 120,000 electron configurations by approximately 4mHa, with a perturbatively corrected energy discrepancy of about 3mHa. This suggests a high quality of the current truncated NQS ansatz close to the HCI wave function. A subsequent selection of roughly 390,000 electron configurations with $\left|\psi_{\theta}\right|>4\times{10}^{-5}$ yields a minor decrease (less than 4mHa) in variational energy, contrasting with a significant drop (about 15mHa) in the variational HCI energy with about 480,000 configurations. This indicates that enhancing NQS accuracy requires a much larger configuration sample $\mathcal{V}$ during optimization.
@@
NQS 156 594 $\mathbf{-0.38978}$ $\mathbf{-0.43753}$
NQS 331 397 $\mathbf{-0.39203}$ $\mathbf{-0.43452}$
@@ Acknowledgements.
Xiang Li: Conceptualization (equal); Data Curation (equal); Formal Analysis (equal); Investigation (equal); Methodology (equal); Project Administration (equal); Software (equal); Supervision (equal); Validation (equal); Visualization (equal); Writing – original draft (equal); Writing – review & editing (equal) Jia-Cheng Huang: Conceptualization (equal); Data Curation (equal); Formal Analysis (equal); Investigation (equal); Methodology (equal); Software (equal); Validation (equal); Visualization (equal); Writing – review & editing (equal) Guang-Ze Zhang: Conceptualization (equal); Data Curation (equal); Formal Analysis (equal); Investigation (equal); Methodology (equal); Software (equal); Validation (equal); Visualization (equal); Writing – review & editing (equal) Hao-En Li: Validation (supporting); Visualization (supporting); Writing – review & editing (equal) Zhu-Ping Shen: Validation (supporting); Visualization (supporting); Writing – review & editing (equal) Chen Zhao: Visualization (supporting); Writing – review & editing (supporting) Jun Li: Funding Acquisition (supporting); Project Administration (supporting); Supervision (supporting); Writing – review & editing (supporting) Han-Shi Hu: Funding Acquisition (equal); Project Administration (equal); Resources (equal); Supervision (equal); Writing – review & editing (equal)
@@
Appendix A Network parameter initialization with molecular orbital energy
We focus on the input biases $a_{i}$ of the RBM, which are uniquely associated with the spin orbitals. If a specific spin orbital is occupied within an electron configuration, denoted as $\sigma_{i}^{k}=1$ in Eq. (2), then the probability $P(D_{k})$ of this configuration includes the term $e^{a_{i}+a_{i}^{*}}$, representing the probability contribution of input bias $a_{i}$. Therefore, we investigate the impact of parameter $a_{i}$ during NQS optimization by evaluating $e^{a_{i}+a_{i}^{*}}$ before and after calculating the ground-state energy of $\rm H_{2}O$ with an O--H bond length of 1.0Å, using the 6-31G basis set.
As depicted in Figure 9, when the parameters are randomly initialized before NQS optimization, the probability contributions of input biases $a_{i}$ are similar, indicating that there is little difference in the probability of electron configurations regardless of which spin orbital the electron occupies. However, a distinctive distribution emerges after optimization, where the occupied spin orbitals with lower energy levels exhibit larger probability contributions, consistent with the fact that electrons prefer to occupy orbitals with lower energy levels.
Figure 9: Graphic representation for the probability contribution of input biases $a_{i}$, in which the horizontal axis denotes the corresponding spin orbitals $\epsilon_{i}$ and is broadened along the vertical direction for clarity. The orbital energy increases from left to right. (a) The randomly initialized input biases before optimization. (b) The randomly initialized input biases after optimization. (c) The input biases initialized using orbital energy with $T=1$ in Eq. (3).
This observation inspired a network parameter initialization strategy using the energy $\epsilon_{i}$ of molecular orbitals, which employs the Boltzmann distribution in Eq. (3) to approximate the optimized probability contribution of each input bias $a_{i}$. As shown in Figure 9, this simple initialization method yields a much better approximation than random initialization.
Appendix B Derivation of the perturbative correction for NQS